OpenStudy (anonymous):

Write an equation of the line containing the given point and perpendicular to the given line. Express the answer in the form of y=mx+b (3,4); 4x+y=5

OpenStudy (anonymous):

4x+y=5 can you tell me the slope of this line??

OpenStudy (anonymous):

I believe its 1 over 4

OpenStudy (dumbcow):

the slope of line is -4 ... 1/4 is slope of perpendicular line now plug in point to get equation: y = mx+ b 4 = (1/4)*3 + b solve for "b" b = 13/4 --> y = 1/4x + 13/4